OpenStudy (anonymous):

On a scale drawing of a park, the sand box is 2 inches wide. The actual sand box is 8 feet wide. What is the scale of the drawing?

OpenStudy (anonymous):

If the drawing is 2 inches wide, and it's actually 8 feet, find a ratio. 2 inches=8 feet. This is a ratio, which can be written as a fraction: 2/8=1/4 Which means 1 inch=4 feet. http://answers.yahoo.com/question/index?qid=20121113065610AAfEs66

OpenStudy (anonymous):

ok thanks! :)

OpenStudy (anonymous):

welcome